A small company is using the unit-of-production method for determining depreciation costs. The original value of the property is $110,000. It is estimated that the company can produce 11,000 units before the equipment will have a salvage or scrap value of zero; that is, the depreciation cost per unit produced is $10. The equipment produces 200 units during the first year, and the production rate is doubled each year for the first 4 years. The production rate obtained in the fourth year is then held constant until the value of the equipment is paid off. What would have been the annual depreciation cost if the straight-line method based on this same time period had been used?


1
Expert's answer
2021-04-07T10:44:41-0400

The straight line for the machine would be calculated as follows:


1.     Cost of the asset: $100,00

2.     Total depreciable cost = cost of the asset - salvage value               

                                        = $100,000 - $0   = $100,000


3.     Useful life of the asset

First Year            200       +

Second Year      400

                             600       +

Third Year          800

                             1400     +

Fourth Year       1600

                             3000     +

Fifth Year           3000

                             6000     +

Sixth Year           3000

                             9000


(11,000 - 9,000)/(11,000)*12 = 0.67 year

 Therefore, the useful life of asset = 6 + 0.67 = 6.67 years


4.     Divide step (2) by (3)

$ 100,000/6.67 = $ 14,992.50


Therefore, it would depreciate at the amount of $ 14,992.50 annually for 6.67 years